geerky42 (geerky42):

The trapezoid shown exists such that BC = 8 cm, AD = 12 cm, BC || EF || AD, and EF and the two diagonals intersect at a common point. Find EF

geerky42 (geerky42):



OpenStudy (campbell_st):

you need to look at the ratios in similar trapeziods.. if EF is parallel to AD and BC then you have 2 trapeziods AEFD and EBCF so looking at the ratios of sides \[\frac{EF}{AD}= \frac{BC}{EF} \] given the known measurements \[\frac{EF}{12} = \frac{8}{EF}\] then after cross multiplying \[EF^2 = 96\] which makes sense as the length of EF will be just less than 10, which is between the 2 known side lengths 8 and 12... I'll leave you to calculate the actual value of EF

geerky42 (geerky42):

@campbell_st Your answer is wrong, I have answer key and the answer is actually 9.6 (nearest tenth). From your answer, I got 9.8...